OpenStudy (anonymous):

If you are implicitly differentiating the first partials, do you differentiate x and view the 's and z's as constants? Then differentiate y and hold x's and z's as constants, etc?

OpenStudy (anonymous):

can you post the actual question? i think i know what you mean, but its hard to answer this question when its not super clear.

OpenStudy (anonymous):

The original question was: Find all the first partials implicitly of (x)lny+(y^2)z+(z^2)=8 I think I figured it out by doing the partial z with respect to x = -F(x)/F(z) and partial z with repect to y = -F(y)/F(z) ....right? maybe?